OpenStudy (blu_seaa):

What are the solutions to the equation 9x2 = 49, rounded to the nearest tenth? A. 2.3, –2.3 B. 5.4, –5.4 C. 7, –7 D. 21, –21 I think it's C.

OpenStudy (standby):

isn't it 7x2=49????

OpenStudy (welshfella):

first divide both sides of the equation by 9 to get the values of x^2 then take the square root

OpenStudy (welshfella):

x^2 = 49/9 x = +/- square root of 49/9

OpenStudy (standby):

no, it should be x2=49 because otherwise we don't see the match cases.

OpenStudy (welshfella):

= sqrt49 / sqrt9 = 7/3

OpenStudy (welshfella):

divide 7 by 3 - what do you get?

OpenStudy (blu_seaa):

2.3

OpenStudy (welshfella):

and ?

OpenStudy (welshfella):

- the negative of 2.3

OpenStudy (blu_seaa):

Oh yeah, thank you.

OpenStudy (welshfella):

sorry i should have said +/- 7/3

OpenStudy (welshfella):

yw

OpenStudy (blu_seaa):

It's fine. c: